OpenStudy (anonymous):

Solve the equation 1 over 2 x + 2 = 22 for x.

OpenStudy (cgreenwade2000):

Is it \[\frac{ 1 }{ 2x+2 } = 22 or \frac{ 1 }{ 2x } + 2 = 22\]